### v1

@@ 5.1 Main findings
Our method of scoring is easier to use than that of Place Pulse (Dubey et al., 2016; Salesses, 2012) or other relative comparison studies (Naik et al., 2014; Ye et al., 2019; Larkin et al., 2022). We collect ratings along an absolute five-point scale, whereas much of the challenge in projects such as Place Pulse was in reconstructing a ranking from a large set of relative comparisons. This is a basic philosophical difference in approach.
@@
The advantage of using absolute scores is that we can already rank photos even with fewer participants and data points. We also argue that with high-level and high-complexity concepts such as ‘walkability’ and ‘safety’ the use of relative comparisons is not necessarily helpful, since there are so many aspects of the imagery to consider simultaneously; past psychological research to justify the use of relative comparisons (Stewart et al., 2005) considered only low-complexity concepts such as ‘line lengths’ or ‘sound tones’. Instead of having a complex process of translating relative comparisons into absolute scores, we keep things simple in our system by showing only one image and enabling participants to give a quick first-impression one-click/swipe answer from a selection of five possible responses. This user interface helps participants quickly and smoothly look through many more images than interfaces with a rating slider (Twedt et al., 2016; Larkin et al., 2022) and avoids any biasing problems created by feeding back preliminary machine learning model results into the user interface (Yao et al., 2019), while still allowing participants to able to enjoy rapid progression through the survey.
@@
We were able to easily attract public participants and give them the flexibility to respond at their convenience because our survey is a smartphone-friendly web app rather than a downloaded software application. We did not pay participants (Twedt et al., 2016; Pearson et al., 2024; Larkin et al., 2022) but rather relied on the attractiveness of the web app as a fun ‘game’ that can be played during downtime (e.g. waiting for the bus).
In the future, we will expand our toolkit by building a framework for generating models of the data collected by our street perception surveys. The state-of-the-art in image classification has moved on considerably in the past five years. Within the past year alone, researchers have published large pre-trained models such as StreetCLIP (Haas et al., 2023), which is designed to perform zero-shot image classification and has been specialised by training on a geo-tagged SVI dataset containing 1.1 million images.
For example, as input to StreetCLIP one could provide a photograph taken from a city street and a set of three descriptive captions such as ‘An urban street’, ‘A suburban road’ and ‘A rural road’. The output of the model would be a probability distribution indicating which one of the input descriptions is the most likely to be suitable for the given image.
@@
StreetCLIP is based on the CLIP text and image association models (Radford et al., 2021). We can take StreetCLIP even further by fine-tuning it on our collected images and ratings by transforming our categories and five-point numerical rating scale into textual image captions using a template such as ‘An image of a street with a [rating text] [category] rating’. For example a walkability rating of 5 becomes ‘An image of a street with a very good walkability rating’, and a safety rating of 2 becomes ‘An image of a street with a bad safety rating’. The resulting image and text models will be able to take an input of an image and a selection of five templated sentences, and then output a probability distribution indicating which of those the most likely sentence to apply to the given image. In this manner, we can extrapolate the perception data by starting from the collected ratings for tens of thousands of images and then obtain modelled ratings for millions of images. One disadvantage of StreetCLIP is that the published model is based on Google SVI, therefore one option we must consider is retraining it based on open or VSVI instead, in addition to our collected dataset.
@@ 5.2 Limitations
Mapillary777mapillary.com/about, as of February 2024 has more than 2 billion images from 190 countries, and especially good coverage in countries like the Netherlands, including almost full panoramic coverage of Amsterdam-area streets. However, although Mapillary has imagery from every continent, the most abundant SVI coverage comes from Europe and North America (Ma et al., 2019).
@@
The absolute scoring system avoids the complexity of the relative comparison system used by Place Pulse (Dubey et al., 2016; Salesses, 2012) but can result in situations where participants might change their minds about images and wish to go back and redo the ratings. We purposefully limited ‘undo’ functionality to a single previous rating to prevent participants from undoing large numbers of ratings. This protects the integrity of the system against unexpected mass cancellation of ratings. Our system collects ‘quick impressions’ with the idea being that participants spend no more than a few seconds on each image; revision or reconsideration of ratings would work against that goal, even if the participant later thinks differently.
@@ C.4 Finding road center-lines
We then apply semantic segmentation (Thisanke et al., 2023), in this case using the Mask2Former model (Cheng et al., 2022) trained on the CityScapes dataset (Cordts et al., 2016), readily available through the HuggingFace transformers library (Wolf et al., 2020). We are only interested in finding the portion of the image that corresponds to ‘road’, therefore Figure 7(a) shows the road pixels highlighted in black, and everything else is grayed out.

### v3

@@ 5.1 Main findings
Our method of scoring is easier to use than that of Place Pulse (Dubey et al., 2016; Salesses, 2012) or other relative comparison studies (Naik et al., 2014; Ye et al., 2019; Larkin et al., 2022). Like Kruse et al. (2021), we collect ratings along an absolute five-point scale, whereas much of the challenge in projects such as Place Pulse was in reconstructing a ranking from a large set of relative comparisons; this is a basic philosophical difference in approach. The advantage of using absolute scores is that the ratings are always linearly ordered. We also argue that with high-level and high-complexity concepts such as ‘walkability’ and ‘safety’ the use of relative comparisons is not necessarily helpful, since there are so many aspects of the imagery to consider simultaneously; past psychological research to justify the use of relative comparisons (Stewart et al., 2005) considered only low-complexity concepts such as ‘line lengths’ or ‘sound tones’. Instead of having a complex process of translating relative comparisons into absolute scores, we keep things simple in our system by showing only one image and enabling participants to give a quick first-impression one-click/swipe answer from a selection of five possible responses. This user interface helps participants quickly and smoothly look through many more images than interfaces with a rating slider (Twedt et al., 2016; Larkin et al., 2022) and avoids any biasing problems created by feeding back preliminary machine learning model results into the user interface (Yao et al., 2019), while still allowing participants to able to enjoy rapid progression through the survey.
@@
We were able to easily attract public participants and give them the flexibility to respond at their convenience because our survey is a smartphone-friendly web app rather than a downloaded software application. We did not pay participants (Twedt et al., 2016; Pearson et al., 2024; Larkin et al., 2022; Kruse et al., 2021) but rather relied on the attractiveness of the web app as a fun ‘game’ that can be played during downtime (e.g. waiting for the bus).
@@
In the future, we will expand our toolkit by building a framework for generating models of the data collected by our street perception surveys. The state-of-the-art in image classification has moved on considerably in the past five years. Within the past year alone, researchers have published large pre-trained models such as StreetCLIP (Haas et al., 2023), which is designed to perform zero-shot image classification and has been specialised by training on a geo-tagged SVI dataset containing 1.1 million images. For example, as input to StreetCLIP one could provide a photograph taken from a city street and a set of three descriptive captions such as ‘An urban street’, ‘A suburban road’ and ‘A rural road’. The output of the model would be a probability distribution indicating which one of the input descriptions is the most likely to be suitable for the given image. StreetCLIP is based on the CLIP text and image association models (Radford et al., 2021). We can take StreetCLIP even further by fine-tuning it on our collected images and ratings by transforming our categories and five-point numerical rating scale into textual image captions using a template such as ‘An image of a street with a [rating text] [category] rating’. For example a walkability rating of 5 becomes ‘An image of a street with a very good walkability rating’, and a safety rating of 2 becomes ‘An image of a street with a bad safety rating’. The resulting image and text models will be able to take an input of an image and a selection of five templated sentences, and then output a probability distribution indicating which of those the most likely sentence to apply to the given image. In this manner, we can extrapolate the perception data by starting from the collected ratings for tens of thousands of images and then obtain modelled ratings for millions of images. One disadvantage of StreetCLIP is that the published model is based on Google SVI, therefore one option we must consider is retraining it based on open or VSVI instead, in addition to our collected dataset.
@@ 5.2 Limitations
Mapillary777mapillary.com/about, as of February 2024 has more than 2 billion images from 190 countries, and especially good coverage in countries like the Netherlands, including almost full panoramic coverage of Amsterdam-area streets. However, although Mapillary has imagery from every continent, the most abundant SVI coverage comes from Europe and North America (Ma et al., 2019). However, commercial SVI is no panacea either: Kim and Jang (2023) analyzed Google SVI coverage of walk commute trajectories in small- to medium-sized cities in the United States and found gaps in nearly half of the routes they investigated; they suggest that researchers consider Mapillary to bolster coverage.
Like with any SVI-based measurements, the outputs of the survey created by this system will depend upon imagery selection choices with regard to directionality, spacing and specific viewing position on the street (e.g. sidewalk-view vs road-centered-view) for each image (Kim et al., 2021; Ki et al., 2023). For the sake of the case study we have made choices specific to the Amsterdam area and the practical considerations we had to make in terms of the level of participation we could reasonably expect. Researchers deploying this toolkit on their own studies will need to consider similar issues specific to their situations.
@@
The absolute scoring system avoids the complexity of the relative comparison system used by Place Pulse (Dubey et al., 2016; Salesses, 2012) but can result in situations where participants might change their minds about images and wish to go back and redo the ratings. We purposefully limited ‘undo’ functionality to a single previous rating to prevent participants from undoing large numbers of ratings. This protects the integrity of the system against unexpected mass cancellation of ratings. Our system collects ‘quick impressions’ with the idea being that participants spend no more than a few seconds on each image; revision or reconsideration of ratings would work against that goal, even if the participant later thinks differently. Relatedly, the absolute scoring system makes it possible to rank images with trivial effort (because absolute scores form a total order), but that ranking could be subject to significant fluctuations when the number of ratings per image is low. This could happen if new ratings for an image are collected that significantly differ from the existing ratings for that image. Therefore, the potential number of participants, and how many images they might be able to rate, should be considered when choosing the number of images to include.
@@ C.4 Finding road center-lines
We then apply semantic segmentation (Thisanke et al., 2023), in this case using the Mask2Former model (Cheng et al., 2022) trained on the CityScapes dataset (Cordts et al., 2016), readily available through the HuggingFace transformers library (Wolf et al., 2020). We chose this off-the-shelf model for no reason other than because it was the most popular and relevant semantic segmentation model available for rapid deployment from Huggingface. We found that it works sufficiently well for the roads we studied in Amsterdam but we do not depend on any particular property of it. Therefore, users preparing their own imagery with our toolkit could substitute another model, if desired. We are only interested in finding the portion of the image that corresponds to ‘road’, therefore Figure 7(a) shows the road pixels highlighted in black, and everything else is grayed out.